Calculate the cash dividends paid out to common stockholders based on the following information: Balance of retained earnings, December 31, 2016 = 388,000,000 Net Income for 2017 44,000,000 Dividends paid to preferred stockholders 1,000,000 Balance of retained earnings, December 31, 2017 372,000,000

(Answer in Millions of dollars, with no 0s or $ sign needed, e.g. xx, or 12)
Answer #1

Ending retained earnings=Beginning retained earnings+Net income-Dividends paid to preferred stockholder-dividends paid out to common stockholders

372,000,000=388,000,000+44,000,000-1,000,000-dividends paid out to common stockholders

dividends paid out to common stockholders=388,000,000+44,000,000-1,000,000-372,000,000

=59,000,000

Hence dividends paid to common stockholders=59 ( in millions of dollars).